OpenStudy (anonymous):

Find the area of a rectangle if the length and width are (3x - 2) inches and (2x+ 4) inches ?

OpenStudy (jdoe0001):

\(\bf \textit{area of a rectangle}=length \times width\\ length = (3x - 2), width=(2x+ 4)\\ \quad \\ \textit{area of a rectangle}=(3x - 2)\times (2x+ 4)\)

OpenStudy (anonymous):

(3x-2) * (2x+4) 6x-8 = Area

OpenStudy (anonymous):

I'm pretty sure that's right.